Ticket Number: HEA-NOC/20060731-3 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20060731 11:52 UTC+1 Problem Start: 20060731 11:47 UTC+1
Ticket Closed: 20060803 15:12 UTC+1 Problem End: 20060731 18:12 UTC+1
Site/line: SVUH, UCD
UCD has no backup link, its primary link is fine.
UCD can not contact St. Vincents and Eccles Street.
We are investigating.
ST.Vicents lost connectivity.
20060731 11:47 LN
20060731 13:05 OMG An Engineer from CWS is en-route to DIT to check the
fibre patching there to Belfield. I have also asked Cieren Perry to reseat
the fibre in the TCD mux.
20060731 13:21 OMG Reseating the fibre in the TCD mux didn't bring up the
connection, so we're still waiting on the contractor to check the fibre in
20060731 17:06 OMG Extensive fibre checks have been done on the fibre to
UCD and it has tested clear, but the link remains down. Have contacted
Lancomms to swap out the CWDM kit in DIT Aungier Street and the remaining
fibre patches to the MUX.
20060731 18:24 OMG Checked the fibre patching on site in DIT and
discovered that the fibre from TCD that should have been plugged into the
west point of the CWDM box was plugged into the 1570nm drop point instead.
I swapped the fibre and all links came back immediately. Will monitor for
20060803 15:11 LN
All connectivity went back. ticket closed.
Time to Fix:
Original problem start was 20060729 09:13. Time to fix 3 1/2 days.
This ticket can be monitored at http://www.hea.net/tickets/20060731-3